There are 120 members in a certain activity club 52 take up dancing 58 joining the debating and 64 learn swimming 14 learn dancing and debating 27 debating and swimming and 32 swimming and dancing if each takes up at least one activity how many take up all three activities?
Answer by ikleyn(52797)

.
We are given 3 sets: D (dancing) of 52 elements;  B (debating) of 58 elements and W (swimming) of 64 elements.


We are also given the number of elements in each in-pair intersections.


We also are told that the union of all three sets contains 120 elements.


In such situation the following formula works


    n(D U B U W) = n(D) + n(B) + n(W) - n(D n B) - n(D n W) - n(B n W) + n(D n B n W).     (*)


Substituting the given data, you get


    120 = 52 + 58 + 64 - 14 - 27 - 32 + n(D n B n W).


In this equation, the only term  n(D n B n W)  is unknown - exactly that under the question.


So, it is easy to get it from the last equation:


    n(D n B n W) = 120 - 52 - 58 - 64 + 14 + 27 + 32 = 19.


Answer.  19 members enjoy all three activities.

Solved.

-----------------

The last step is to prove the formula (*).

    It is totally clear to you why I add the first three addends in the formula (*).


    But when I added them, I counted twice every term in each in-pair intersection.


    Therefore, I subtracted the number of terms in each in-pair intersection.


    Next, when I added three first addends, I counted thrice each term in the triple intersection;

    and when I subtracted in-pair intersections, I canceled these terms thrice.

    Therefore, I must add the number of terms in the triple intersection one more time to restore the balance.

Thus the formula  (*)  is proved   //   and the solution is completed  (! !)

------------------

See the lessons
    - Counting elements in sub-sets of a given finite set
    - Advanced problems on counting elements in sub-sets of a given finite set
    - Challenging problems on counting elements in subsets of a given finite set
in this site.